# On computing and the complexity of computing higher-order $U$-statistics, exactly

### v1

@@ 4.2 Application 2: Motif Counts in Random Graph Models and Networks
where the kernel function $h_{\mathsf{r}}$ encodes the motif structure, and $|\mathrm{Aut}(\mathsf{r})|$ denotes the number of automorphisms of motif $\mathsf{r}$, which counts the number of symmetric repetitions.
To be concrete, letting $B\equiv 1-A$, the corresponding $U$-statistics for motif counts for specific $3$-vertex motifs and $4$-vertex motifs (see Figure 2 for which motifs $\mathsf{r}_{1}$ to $\mathsf{r}_{8}$ encode) read as follows:
We compare our package u-stats with igraph (csardi2006igraph) on exactly computing all 3-vertex and all 4-vertex motif counts. igraph computes motif counts by using a highly optimized algorithm based on wernicke2006fanmod.
We generate random graphs using the classical Erdös-Rényi model $G(n,p)$, where each of the $\binom{n}{2}$ possible edges between $n$ vertices is present independently with probability $p$. We report the runtime comparisons in Table 4 and Table 5. All experiments were conducted on a server equipped with Intel Xeon Platinum 8358 CPUs (2.6GHz, 64 cores) and memory of 512 GB. Our implementation makes full use of all 64 CPU cores via multiprocessing, while igraph runs in a single-threaded setting, due to its current implementation limitation.
To quantify the performance, we consider two metrics:
  • Speedup: the ratio of igraph runtime to the runtime of u-stats with parallel computation;
  • Single-core speedup: a very rough estimate of per-core speedup, computed as the speedup divided by the number of cores (i.e. $64$).
As shown in Table 4 and Table 5, igraph has better runtime performance than u-stats when $p$ is small, corresponding to a sparse graph. Because u-stats is designed as a general purpose package for computing $U$-statistics, we do not further take into account the potential sparsity of the adjacency matrix. When $p$ increases (i.e., as the graph becomes denser), the runtime of igraph grows considerably even when $p$ is as low as $0.1$. Yet u-stats still maintains roughly the same runtime throughout, which is much faster than that of igraph when $p$ grows.
(a) $\mathsf{r}_{1}$: V-shape
(b) $\mathsf{r}_{2}$: Triangle
(c) $\mathsf{r}_{3}$ : 3-star
(d) $\mathsf{r}_{4}$ : Fork
(e) $\mathsf{r}_{5}$ : Tailed triangle
(f) $\mathsf{r}_{6}$: Square
(g) $\mathsf{r}_{7}$: House
(h) $\mathsf{r}_{8}$ : 4-clique
@@
Table 4: Runtime comparison of exact all 33-vertex motif counts using u-stats and igraph on Erdös-Rényi graphs $G(n,p)$ with $n=5623$. Experiments were run on Intel Xeon ICX Platinum 8358 CPUs (2.6GHz, 64 total cores) with memory of 512 GB. “Speedup ratio” compares igraph to our parallel method; “single core speedup” assumes single-core execution.
Edge
Prob. $p$
u-stats
Time (s)
igraph
Time (s)
Speedup
Ratio
Single-Core
Speedup
0.001 0.617 0.025 0.040 0.001
0.005 0.649 0.353 0.544 0.008
0.010 0.692 1.795 2.595 0.041
0.020 0.783 10.594 13.529 0.211
0.050 1.038 136.272 131.288 2.051
0.080 1.298 514.702 396.415 6.194
0.100 1.453 960.299 660.690 10.323
0.150 1.864 3003.490 1611.389 25.178
0.200 2.275 6725.515 2955.656 46.182
Table 5: Runtime comparison of exact all 44-vertex motif counts using u-stats and igraph on Erdös-Rényi graphs $G(n,p)$ with $n=2000$. Experiments were run on Intel Xeon ICX Platinum 8358 CPUs (2.6GHz, 64 total cores) with memory of 512 GB. “Speedup ratio” compares igraph to our parallel method; “single core speedup” assumes single-core execution.
Edge
Prob. $p$
u-stats
Time (s)
igraph
Time (s)
Speedup
Ratio
Single-Core
Speedup
0.001 70.959 0.004 0. 0.
0.005 71.947 0.168 0.002 0.
0.010 68.463 1.453 0.021 0.
0.020 65.634 15.573 0.237 0.004
0.050 65.786 394.532 5.997 0.094
0.080 66.866 2371.373 35.465 0.554
0.100 68.219 5396.057 79.099 1.236
0.150 65.969 23003.041 348.692 5.448
0.200 65.259 61796.760 946.940 14.796
4.3 Application 3: Distance Covariance Measures for Independence Testing
@@
In the last example, we revisit the problem considered in Section 5.1 of shao2025u, which concerns the problem of computing the squared distance covariance $\mathsf{dCov}^{2}$, a recently proposed dependence measure between two random vectors (szekely2007measuring; yao2018testing). In particular, let $O_{i}\coloneqq(X_{i},Y_{i}),i\in[n]$ be the observed sample. $\mathsf{dCov}^{2}(X,Y)$ can then be written as a $4$-th order $U$-statistic
@@
where $h_{1}(O_{i},O_{j})\coloneqq\|X_{i}-X_{j}\|_{2}$ and $h_{2}(O_{i},O_{j})\coloneqq\|Y_{i}-Y_{j}\|_{2}$ for $i,j\in[n]$. Here $\|\cdot\|_{2}$ denotes the vector $\ell_{2}$-norm.
Quoting from Table 5 of shao2025u, it takes “8099.73 seconds” to compute $\mathsf{dCov}^{2}$ with $n=138$ among 11 random vectors. Each vector corresponds to a different stock sector—such as Health Care or Industrials—and its components represent the monthly log-returns of individual stocks within that sector, using historical S&P 500 data. To bypass such a high computational burden, shao2025u resort to computing a randomized incomplete $U$-statistic instead. In the same Table 5, shao2025u demonstrated that there is indeed a significant speedup when using randomized incomplete $U$-statistics. It is noteworthy that another motivation of shao2025u to use randomized incomplete $U$-statistic is due to its asymptotic normality (chen2019randomized) even when the original complete $U$-statistic has a Gaussian chaos limiting distribution. But we will not further discuss this issue as it is not directly relevant to the main theme of our paper.
In Table 6, we display the average runtime over 10 runs of exactly computing $\mathsf{dCov}^{2}$ using our package u-stats, or computing $\mathsf{dCov}^{2}$ exactly or approximately with the randomized incomplete $U$-statistic using the MATLAB code attached in the supplementary materials of shao2025u. We rerun the code of shao2025u together with u-stats on the same computing platform to ensure a fair comparison. In particular, the algorithm proposed in shao2025u will randomly choose $O(n^{\alpha})$ permutations, with $\alpha>0$ a tuning parameter determined by the available computation budget. By choosing $\alpha<4$, the running time will be reduced compared to $O(n^{4})$. We vary the tuning parameter $\alpha$ within the set $\{1.5,2,2.5\}$. Our package u-stats again exhibits quite impressive runtime performance: u-stats without parallel computation takes slightly shorter time than the randomized incomplete $U$-statistic when $\alpha=2$, whereas u-stats with parallel computation takes less than half of the time as the case when $\alpha=1.5$!
@@
No Parallel
Parallel
Randomized
$\alpha=1.5$
Randomized
$\alpha=2.0$
Randomized
$\alpha=2.5$
Complete
@@ 5 Concluding Remarks
In this paper, we have conducted a delicate and systematic study on computing and the complexity of computing higher-order $U$-statistics (and also $V$-statistics), exactly. Explicit connections between the complexity of computing higher-order UU/$V$-statistics and concepts from tensor computation and graph theory are established. An accompanying Python package u-stats is provided to help practitioners compute UU/$V$-statistics more efficiently in applications. Although Python becomes more and more popular among statisticians (dai2023rehline), an R interface of u-stats will also be available in due course. It will also be interesting to test u-stats on other problems involving $U$-statistics (qiu2024neural; shen2025engression).
One limitation of our work is that we prioritize time complexity over space complexity. As a result, when we need to store a very high-order tensor, the sample size $n$ cannot be too large – this is why we have only evaluated the performance of u-stats on computing HOIF estimators with order $m\leq 7$ when $n=10000$. An important future direction is thus to explore how to balance the time-complexity and space-complexity when the sample size $n$ is large.
Finally, we conclude our paper with an open problem that we ourselves find intriguing. As we indicated in Remark 7 at the end of Section 3, the treewidth of the decomposition graph associated with the Einsum notation of the underlying $V$-statistic relates to its time complexity. But the treewidth of a graph is not easy to compute either analytically (harvey2018treewidth) or numerically (arnborg1987complexity). It will be interesting to get a sharp bound on treewidth using the number of vertices or edges of the graph. As mentioned in Remark 7, kneis2009bound has proved that the treewidth is bounded at most linearly by the edge size, or more precisely bounded by $e/5.769+O(\log n)$ where $n\coloneqq|V|$ denotes the number of vertices in $G$. Yet we do not know if this bound can be further improved. Below we provide the maximum tree-width possible for edge size no greater than $15$, to hopefully serve as a starting point of fully addressing this problem.
Observation 1 (Treewidth of Graphs with a Given Number of Edges).
Let
@@
The proof can be found in Appendix C.1. However, extending the above result to $e>15$ is currently beyond our reach.
Appendix A Proof of Proposition 1
@@
Definition 13 (Einsum Notation Sequence under a Particular Summation Ordering).
Let ${\mathcal{A}}=(A_{1},\cdots,A_{K})$ be an Einsum notation (recall from Definition 4) with no output of. Let $I=[m]$ be the index set of $\mathcal{A}$} and $\sigma$ be a permutation over $I$. The Einsum notation sequence of $\mathcal{N}$} under $\sigma$ is the sequence ${\mathcal{A}}^{\sigma}_{0},{\mathcal{A}}^{\sigma}_{1},\ldots,{\mathcal{A}}^{\sigma}_{m}$ defined recursively as follows:
@@
Lemma 3 (The Einsum Notation Sequence, Vertex Elimination Correspondence).
@@
We immediately have |B|=d1|B|=d-1, which implies that the result of $\mathsf{Einsum}({\mathcal{T}},{\mathcal{N}})$ has $n^{d-1}$ entries. For each entry, as one index is summed over, it takes n1n-1 summation operations and $(|{\mathcal{A}}|-1)n$ multiplication operations. Therefore, the total time complexity of computing $\mathsf{Einsum}({\mathcal{T}},{\mathcal{N}})$ is $\big{(}(n-1)n^{d-1}+(|{\mathcal{A}}|-1)n^{d}\big{)}$, which is $\Theta(|{\mathcal{A}}|n^{d})$. ∎
[…]
@@
We first show that there is an algorithm such that the time complexity of computing $\mathsf{Einsum}({\mathcal{T}},{\mathcal{A}})$ is $O(n^{C_{\sigma}+1})$ where
@@
  • Let $B^{\sigma,\mathrm{out}}_{i}$ be a permutation of the set $\bigcup_{\begin{subarray}{c}A\in{\mathcal{A}}^{\sigma}_{i-1}\\ \sigma[i]\in A\end{subarray}}(\mathsf{set}(A)\setminus\{\sigma[i]\})$ as the output tuple and $T^{\sigma,\mathrm{out}}_{i}\coloneqq\mathsf{Einsum}({\mathcal{T}}^{\sigma,\mathrm{in}}_{i},({\mathcal{A}}^{\sigma,\mathrm{in}}_{i},B^{\sigma,\mathrm{out}}_{i}))$;
@@
$\mathsf{Einsum}({\mathcal{T}}^{\sigma}_{i},{\mathcal{A}}^{\sigma}_{i})\equiv\mathsf{Einsum}({\mathcal{T}}^{\sigma}_{j},{\mathcal{A}}^{\sigma}_{j})\equiv\mathsf{Einsum}({\mathcal{T}},{\mathcal{A}}).$
Thus the time complexity of computing $\mathsf{Einsum}({\mathcal{T}},{\mathcal{A}})$ just comes from transferring from one step to the next step, which is $\mathsf{Einsum}({\mathcal{T}}^{\sigma,\mathrm{in}}_{i},({\mathcal{A}}^{\sigma,\mathrm{in}}_{i},B^{\sigma,\mathrm{out}}_{i}))$. Following Lemma 4, we obtain that the time complexity of computing $\mathsf{Einsum}({\mathcal{T}}^{\sigma,\mathrm{in}}_{i},({\mathcal{A}}^{\sigma,\mathrm{in}}_{i},B^{\sigma,\mathrm{out}}_{i}))$ is $\Theta(|{\mathcal{A}}^{\sigma,\mathrm{in}}_{i}|n^{\deg(x,G_{{\mathcal{A}}^{\sigma}_{i-1}})})$, which is also $\Theta(|{\mathcal{A}}|n^{\deg(\sigma[i],G_{{\mathcal{A}}^{\sigma}_{i-1}})})$ as $|{\mathcal{A}}^{\sigma,\mathrm{in}}_{i}|\leq|{\mathcal{A}}|$. By taking maximum over $i$, the overall time complexity is thus $O(|{\mathcal{A}}|n^{C_{\sigma}+1})$.
@@
and by finding the permutation $\sigma^{\ast}$ minimizing $C_{\sigma}$ and recalling Definition 7, we can conclude that there exists an algorithm such that the time complexity of $\mathsf{Einsum}({\mathcal{T}},{\mathcal{A}})$ is $O(|{\mathcal{A}}|n^{\mathsf{tw}(G_{\mathcal{A}})+1})$, which completes the proof. ∎
Appendix B Proof of Lemma 1
@@
${\mathcal{U}}\left(n,\pi\right)\coloneqq\left\{(s_{1},s_{2},\ldots,s_{m})\in[n]^{m}\mid\text{where for any $i,j\in[m]$},i,j\in Q\text{ for some $Q\in\pi$}\iff s_{i}=s_{j}\right\}.$
@@
Definition 16 (Restricted $U$-statistics).
@@
The proof of Lemma 5 can be found in Appendix B.2.
B.1 Preparatory Materials
@@
Lemma 6 (Decomposition of a $V$-statistic into $U$-statistics).
@@
Definition 17 (Möbius Function, (stanley2011enumerative)).
@@
Lemma 7 (Möbius Function on Partition Lattices, (stanley2011enumerative)).

### v2

@@ 4.2 Application 2: Motif Counts in Networks
[…]
We compare our package u-stats with several state-of-the-art systems for motif counting, including the widely used graph analysis library igraph (csardi2006igraph), the parallel motif counting system Peregrine (jamshidi2020peregrine), and NVIDIA’s GPU-based library cuGraph (fender2022rapids). These systems represent different computational paradigms. igraph relies on highly optimized enumeration algorithms derived from wernicke2006fanmod and runs in a single-threaded setting. Peregrine implements parallel, motif-specific optimizations for fast CPU-based counting. cuGraph provides specialized GPU kernels, but currently supports only a limited set of motifs (e.g., triangles). In contrast, u-stats adopts a unified formulation based on $U$-statistics and Einsum, without requiring implementations specific to motif counting.
[…]
The results reveal distinct performance regimes across methods. On sparse graphs (small $p$), Peregrine and igraph outperform u-stats due to their ability to exploit sparsity through enumeration-based strategies. However, as the graph becomes denser, their runtime increases significantly, with igraph failing to complete within the time limit even at moderate densities. In contrast, u-stats maintains nearly constant runtime across all values of $p$, resulting in substantial performance advantages in dense regimes. The results for 4-vertex motifs in Table 7 exhibit similar trends. On GPUs, cuGraph achieves faster runtimes on sparse graphs due to specialized kernels, but its performance degrades with increasing density and eventually runs out of memory, whereas u-stats remains stable without requiring motif-specific implementations.
@@
Table 2: Runtime comparison of exact all $3$-vertex motif counts using u-stats, Peregrine and igraph on Erdős–Rényi graphs $G(n,p)$ with $n=20000$. Experiments were run on Intel Xeon Scalable Cascade Lake 6248 CPUs (2.5GHz, 40 total cores) with memory of 192 GB. “OOT” denotes instances that exceeded the time limit of 1800 seconds.
Edge Prob. $p$ u-stats Time (s) Peregrine Time (s) igraph Time (s)
0.0005 10.3478 0.0703 0.1915
0.0010 10.4452 0.0740 0.7259
0.0050 10.6196 0.0780 40.7373
0.0100 10.3616 0.1190 289.5082
0.0500 9.8357 1.1782 OOT
0.2000 10.6010 17.1063 OOT
0.4000 10.8445 67.6727 OOT
0.8000 8.8623 158.7377 OOT
4.3 Application 3: Distance Covariance Measures
@@
In the last example, we revisit the problem considered in Section 5.1 of shao2025u, which concerns the problem of computing the squared distance covariance $\mathsf{dCov}^{2}$, a recently proposed dependence measure between two random vectors (szekely2007measuring; yao2018testing). In particular, let $O_{i}\coloneqq(X_{i},Y_{i}),i\in[n]$ be the observed sample. $\mathsf{dCov}^{2}(X,Y)$ can be written as a $4$-th order $U$-statistic: Denoting $\|\cdot\|_{2}$ as the vector $\ell_{2}$-norm, then
@@
where $h_{1}(O_{i},O_{j})\coloneqq\|X_{i}-X_{j}\|_{2}$ and $h_{2}(O_{i},O_{j})\coloneqq\|Y_{i}-Y_{j}\|_{2}$ for $i,j\in[n]$.
Quoting from Table 5 of shao2025u, it takes “8099.73 seconds” to compute $\mathsf{dCov}^{2}$ among 11 random vectors, together with its confidence interval, when $n=138$. Each vector corresponds to a different stock sector—such as Health Care or Industrials—and its components represent the monthly log-returns of individual stocks within that sector, using historical S&P 500 data. shao2025u resort to computing a randomized incomplete $U$-statistic instead. In the same Table 5, shao2025u demonstrated that there is indeed a significant speedup when using randomized incomplete $U$-statistics.
In Table 3, we display the average runtime over 10 runs of exactly computing $\mathsf{dCov}^{2}$ using our package u-stats, or computing $\mathsf{dCov}^{2}$ exactly or approximately with the randomized incomplete $U$-statistic using the MATLAB code attached in the supplementary materials of shao2025u. We rerun the code of shao2025u together with u-stats on the same computing platform to ensure a fair comparison and we only focus on point estimation. In particular, the algorithm proposed in shao2025u will randomly choose $O(n^{\alpha})$ permutations, with $\alpha>0$ a tuning parameter determined by the available computation budget. By choosing $\alpha<4$, the running time will be reduced compared to $O(n^{4})$. We vary the tuning parameter $\alpha$ within the set $\{1.5,2,2.5\}$. Our package u-stats again exhibits quite impressive runtime performance: u-stats without parallel computation takes slightly shorter time than the randomized incomplete $U$-statistic when $\alpha=2$, whereas u-stats with parallel computation takes less than half of the time of the case when $\alpha=1.5$!
@@
No Parallel Parallel Randomized $\alpha=1.5$ Randomized $\alpha=2.0$ Randomized $\alpha=2.5$ Complete
@@ 5 Concluding Remarks
In this paper, we have conducted a delicate and systematic study on exactly computing and the complexity of exactly computing higher-order $U$-statistics (and also $V$-statistics). Explicit connections are established between the complexity of computing higher-order UU/$V$-statistics and concepts from tensor computation and graph theory. An accompanying Python package u-stats and its R interface are provided to help practitioners compute UU/$V$-statistics more efficiently in applications. It will also be interesting to test u-stats on other problems involving $U$-statistics (shen2025engression).
One limitation of our work is that we prioritize time complexity over space complexity. As a result, when we need to store a very high-order tensor, the sample size $n$ cannot be too large; this is why we have only evaluated the performance of u-stats on computing HOIF estimators of order $m\leq 7$ when $n=10000$. Therefore, an important future direction is to explore how to balance time and space complexities when both $m$ and $n$ are even larger.
SUPPLEMENTARY MATERIAL
@@
In this section, we provide two additional examples to illustrate the definitions of the Einsum notation and the Einsum operation, supplementing Example 1.
Example 2.
We take matrix multiplication as an example to illustrate the Einsum operation just defined. Let $T_{1},T_{2}$ be two input matrices, which are also 2nd-order tensors. Therefore $\mathcal{T}=(T_{1},T_{2})$. Suppose we are interested in multiplying $T_{1}$ and $T_{2}$, denoted as $T^{\rm out}$: for any $i,j\in[n]$, $T^{\rm out}(i,j)=\sum\limits_{k=1}^{n}T_{1}(i,k)T_{2}(k,j)$. The matrix multiplication corresponds to the Einsum operation with Einsum notation $\mathcal{N}=(\mathcal{A};B)$, with input tuple list $\mathcal{A}=((1,2),(2,3))$ and output tuple $B=(1,3)$ and $|B|=2$. The index set is simply $I=[3]$. This means that the operation sums over the dimensions of $T_{1}$ and $T_{2}$ corresponding to index $2$ and reserves indices 11 and $3$. $T^{\rm out}(\alpha)$ for any $\alpha=(i,j)\in[n]^{2}$ can be written in the Einsum operation form $\textsc{Einsum}(\mathcal{T},\mathcal{N})$ because by definition:
$\textsc{Einsum}(\mathcal{T},\mathcal{N})$ $\equiv\sum_{\begin{subarray}{c}\alpha^{\prime}\in[n]^{3}\\ \alpha^{\prime}[1,3]=\alpha\end{subarray}}\prod_{k=1}^{2}T_{k}(\alpha^{\prime}[A_{k}])=\sum_{\begin{subarray}{c}\alpha^{\prime}\in[n]^{3}\\ \alpha^{\prime}[1,3]=\alpha\end{subarray}}T_{1}(\alpha^{\prime}[A_{1}])T_{2}(\alpha^{\prime}[A_{2}])$
$=\sum_{\begin{subarray}{c}\alpha^{\prime}\in[n]^{3}\\ \alpha^{\prime}[1,3]=\alpha\end{subarray}}T_{1}(\alpha^{\prime}[1,2])T_{2}(\alpha^{\prime}[2,3])=\sum_{\begin{subarray}{c}\alpha^{\prime}\in[n]^{3}\\ \alpha^{\prime}[1,3]=\alpha\end{subarray}}T_{1}(\alpha^{\prime}[1],\alpha^{\prime}[2])T_{2}(\alpha^{\prime}[2],\alpha^{\prime}[3])$
$=\sum_{\alpha^{\prime}[2]=1}^{n}T_{1}(\alpha[1],\alpha^{\prime}[2])T_{2}(\alpha^{\prime}[2],\alpha[3])=\sum_{\alpha^{\prime}[2]=1}^{n}T_{1}(i,\alpha^{\prime}[2])T_{2}(\alpha^{\prime}[2],j),$
so the claim holds by identifying $\alpha^{\prime}[2]=k$.
Example 3.
$T^{\rm out}(i,j,k)$ is in fact a partial trace, a common operation taken in quantum mechanics. The corresponding Einsum notation is $\mathcal{N}=(\mathcal{A};B)$ with $\mathcal{A}=(A_{1}=(1,2,3),A_{2}=(2,4,5),A_{3}=(3,6))$ with $B=(1,5,6)$ and the index set $I=[6]$. $T^{\rm out}(\alpha)$ for any $\alpha=(i,j,k)\in[n]^{3}$ can be written in the Einsum operation form $\textsc{Einsum}(\mathcal{T},\mathcal{N})$ because by definition:
$\textsc{Einsum}(\mathcal{T},\mathcal{N})$ $\equiv\sum_{\begin{subarray}{c}\alpha^{\prime}\in[n]^{4}\\ \alpha^{\prime}[1,5,6]=\alpha\end{subarray}}T_{1}(\alpha^{\prime}[A_{1}])T_{2}(\alpha^{\prime}[A_{2}])T_{3}(\alpha^{\prime}[A_{3}])$
$=\sum_{\begin{subarray}{c}\alpha^{\prime}\in[n]^{4}\\ \alpha^{\prime}[1,5,6]=\alpha\end{subarray}}T_{1}(\alpha^{\prime}[1,2,3])T_{2}(\alpha^{\prime}[2,4,5])T_{3}(\alpha^{\prime}[3,6])$
$=\sum_{\alpha^{\prime}[2]=1}^{n}\sum_{\alpha^{\prime}[3]=1}^{n}\sum_{\alpha^{\prime}[4]=1}^{n}T_{1}(\alpha[1],\alpha^{\prime}[2],\alpha^{\prime}[3])T_{2}(\alpha^{\prime}[2],\alpha^{\prime}[4],\alpha[5])T_{3}(\alpha^{\prime}[3],\alpha[6]).$
Thus the claim holds by identifying $\alpha^{\prime}[2]=a$, $\alpha^{\prime}[3]=b$, and $\alpha^{\prime}[4]=c$.
[…]
In he2021asymptotically; lai2023block, higher-order $U$-statistics are employed to test diagonal and block-diagonal structures of high-dimensional covariance matrices. A detailed computational analysis, including an iterative algorithm and complexity bounds, is provided in he2021asymptotically.
Consider the sample $\bm{X}=(X_{i})_{i=1}^{n}$ with $X_{i}\in\mathbb{R}^{p}$. The $U$-statistic used to construct the test (omitting the normalizing constant) takes the form
$\mathcal{U}(a)=\sum_{(j_{1},j_{2})\in\mathbf{I}}\sum_{\bar{i}_{2a}\in\mathsf{Perm}([n],2a)}\prod_{k=1}^{a}(X_{i_{2k-1},j_{1}}X_{i_{2k-1},j_{2}}-X_{i_{2k-1},j_{1}}X_{i_{2k},j_{2}}),$ (6)
where $\mathbf{I}\subseteq[p]^{2}$ depends on the specific testing goal. For diagonal testing, he2021asymptotically uses $\mathbf{I}=[p]^{2}\setminus\bigcup_{j=1}^{p}\{(j,j)\}$.
When the mean of each $X_{i}$ is known to be zero, the statistic simplifies to
$\widetilde{\mathcal{U}}(a)=\sum_{(j_{1},j_{2})\in\mathbf{I}}\sum_{\bar{i}_{a}\in\mathsf{Perm}([n],a)}\prod_{k=1}^{a}(X_{i_{k},j_{1}}X_{i_{k},j_{2}}).$
For this simplified form, he2021asymptotically presents an iterative algorithm (which is a special case of our algorithm) and establishes a computational complexity of $O(p^{2}n)$ independent of the order $a$.
In terms of the original test statistic $\mathcal{U}(a)$ in (6), however, he2021asymptotically only provides an algorithm and complexity analysis for $a\leq 2$. They remark that:
When $a\geq 3$, a similar iterative method can be applied, but a closed-form expression may be difficult to derive directly.
The difficulty in obtaining a closed-form expression stems precisely from the decomposition challenge addressed in Lemma 1 of our paper, which expresses a general $U$-statistic as a linear combination of $V$-statistics. To circumvent this issue, he2021asymptotically instead considers a modified test statistic $\mathcal{U}_{c}(a)$ resembling $\widetilde{\mathcal{U}}(a)$ and shows that it exhibits asymptotic properties similar to $\mathcal{U}(a)$.
Our algorithm u-stats can directly compute $\mathcal{U}(a)$, and our complexity analysis developed in Section 3.3 yields the same runtime complexity $O(p^{2}n)$. To see this, we first expand the product using the binomial theorem:
$\mathcal{U}(a)$ $=\sum_{(j_{1},j_{2})\in\mathbf{I}}\sum_{\bar{i}_{2a}\in\mathsf{Perm}([n],2a)}\sum_{S\subseteq[a]}(-1)^{|S|}\prod_{k\notin S}(X_{i_{2k-1},j_{1}}X_{i_{2k-1},j_{2}})\prod_{k\in S}(X_{i_{2k-1},j_{1}}X_{i_{2k},j_{2}})$
$=\sum_{(j_{1},j_{2})\in\mathbf{I}}\sum_{S\subseteq[a]}(-1)^{|S|}\,\mathbb{U}(h_{S,j_{1},j_{2}},\bm{X}),$
where
$h_{S,j_{1},j_{2}}(\bar{X}_{2a})=\prod_{k\notin S}(X_{i_{2k-1},j_{1}}X_{i_{2k-1},j_{2}})\prod_{k\in S}(X_{i_{2k-1},j_{1}})\prod_{k\in S}(X_{i_{2k},j_{2}}).$
Thus, $h_{S,j_{1},j_{2}}$ admits a MD structure (Definition 9) with
$\mathcal{H}_{S,j_{1},j_{2}}$ $=\bigl(\underbrace{h_{1}(X_{i})=X_{i,j_{1}}X_{i,j_{2}}}_{\text{repeated }a-|S|\text{ times}},\ \underbrace{h_{2}(X_{i})=X_{i,j_{1}}}_{\text{repeated }|S|\text{ times}},\ \underbrace{h_{3}(X_{i})=X_{i,j_{2}}}_{\text{repeated }|S|\text{ times}}\bigr),$
$\mathcal{A}_{S}$ $=\bigl((2k-1)_{i\notin S},(2k-1)_{i\in S},(2k)_{i\in S}\bigr).$
The corresponding decomposition graph $G_{\mathcal{A}_{S}}$ (Figure 2) consists solely of isolated vertices (i.e., it is edgeless). Consequently, all quotient graphs are also edgeless and have treewidth 0.
By Corollary 2, each $\mathbb{U}(h_{S,j_{1},j_{2}},\bm{X})$ can be computed in $O(n)$ time for fixed $S$ and $(j_{1},j_{2})$. Since $|\mathbf{I}|=O(p^{2})$ and there are $2^{a}$ possible subsets $S\subseteq[a]$, the total complexity of computing $\mathcal{U}(a)$ is $O(2^{a}p^{2}n)$. In the regime where $a$ is small relative to $n$ and $p$ (the setting considered in he2021asymptotically), this reduces to $O(p^{2}n)$.
[…]
Figure 2: The decomposition graph $G_{\mathcal{A}_{S}}$, which consists of $a+|S|$ isolated vertices.
[…]
@@
Definition 13 (Einsum Notation Sequence).
Let $\mathcal{A}=(A_{1},\cdots,A_{K})$ be an Einsum notation (recall from Definition 4) with no output. Let $I=[m]$ be the index set of $\mathcal{A}$ and $\sigma$ be a permutation over $I$. The Einsum notation sequence of $\mathcal{A}$ under $\sigma$ is the sequence $\mathcal{A}^{\sigma}_{0},\mathcal{A}^{\sigma}_{1},\ldots,\mathcal{A}^{\sigma}_{m}$ defined recursively as follows:
@@
Lemma 3.
@@
We immediately have |B|=d1|B|=d-1, which implies that the result of $\textsc{Einsum}(\mathcal{T},\mathcal{N})$ has $n^{d-1}$ entries. For each entry, as one index is summed over, it takes n1n-1 summation operations and $(|\mathcal{A}|-1)n$ multiplication operations. Therefore, the total time complexity of computing $\textsc{Einsum}(\mathcal{T},\mathcal{N})$ is $\big((n-1)n^{d-1}+(|\mathcal{A}|-1)n^{d}\big)$, which is $\Theta(|\mathcal{A}|n^{d})$. ∎
We need to introduce another notation before presenting the final proof: given two lists of tuples $\mathcal{A}$ and $\mathcal{B}$, we write $\mathcal{A}\sqsubseteq\mathcal{B}$ to mean that $\mathcal{A}$ is a sublist of $\mathcal{B}$.
@@
We first show that there is an algorithm such that the time complexity of computing $\textsc{Einsum}(\mathcal{T},\mathcal{A})$ is $O(n^{C_{\sigma}+1})$ where
@@
  • Let $B^{\sigma,\mathrm{out}}_{i}$ be a permutation of the set $\bigcup_{\begin{subarray}{c}A\in\mathcal{A}^{\sigma}_{i-1}\\ \sigma[i]\in A\end{subarray}}(\mathsf{set}(A)\setminus\{\sigma[i]\})$ as the output tuple and $T^{\sigma,\mathrm{out}}_{i}\coloneqq\textsc{Einsum}(\mathcal{T}^{\sigma,\mathrm{in}}_{i},(\mathcal{A}^{\sigma,\mathrm{in}}_{i},B^{\sigma,\mathrm{out}}_{i}))$;
@@
$\textsc{Einsum}(\mathcal{T}^{\sigma}_{i},\mathcal{A}^{\sigma}_{i})\equiv\textsc{Einsum}(\mathcal{T}^{\sigma}_{j},\mathcal{A}^{\sigma}_{j})\equiv\textsc{Einsum}(\mathcal{T},\mathcal{A}).$
Thus the time complexity of computing $\textsc{Einsum}(\mathcal{T},\mathcal{A})$ just comes from transferring from one step to the next step, which is $\textsc{Einsum}(\mathcal{T}^{\sigma,\mathrm{in}}_{i},(\mathcal{A}^{\sigma,\mathrm{in}}_{i},B^{\sigma,\mathrm{out}}_{i}))$. Following Lemma 4, we obtain that the time complexity of computing $\textsc{Einsum}(\mathcal{T}^{\sigma,\mathrm{in}}_{i},(\mathcal{A}^{\sigma,\mathrm{in}}_{i},B^{\sigma,\mathrm{out}}_{i}))$ is $\Theta(|\mathcal{A}^{\sigma,\mathrm{in}}_{i}|n^{\deg(x,G_{\mathcal{A}^{\sigma}_{i-1}})})$, which is also $\Theta(|\mathcal{A}|n^{\deg(\sigma[i],G_{\mathcal{A}^{\sigma}_{i-1}})})$ as $|\mathcal{A}^{\sigma,\mathrm{in}}_{i}|\leq|\mathcal{A}|$. By taking maximum over $i$, the overall time complexity is thus $O(|\mathcal{A}|n^{C_{\sigma}+1})$.
@@
and by finding the permutation $\sigma^{\ast}$ minimizing $C_{\sigma}$ and recalling Definition 7, we can conclude that there exists an algorithm such that the time complexity of $\textsc{Einsum}(\mathcal{T},\mathcal{A})$ is $O(|\mathcal{A}|n^{\mathsf{tw}(G_{\mathcal{A}})+1})$, which completes the proof. ∎
[…]
@@
$\mathcal{U}\left(n,\pi\right)\coloneqq\left\{\overline{s}_{m}\in[n]^{m}\mid\text{where for any $i,j\in[m]$},i,j\in Q\text{ for some $Q\in\pi$}\iff s_{i}=s_{j}\right\}.$
@@
Definition 16 (Restricted $U$-Statistic).
@@
[…]
@@
Lemma 6 (Decomposition of $V$-Statistic into $U$-Statistics).
@@
Definition 17 (Möbius Function, stanley2011enumerative).
@@
Lemma 7 (Möbius Function on Partition Lattices, stanley2011enumerative).